OpenStudy (anonymous):

Find y' for this equation (1-x) (dy/dx) - xy = 1 ??? SOS please.

OpenStudy (apoorvk):

Lol SOS y' is nothing but dy/dx itself. So, rearrange the whole equation above to write dy/dx implicitly.

OpenStudy (anonymous):

do you mean find y?

OpenStudy (anonymous):

it's a linear DE.

OpenStudy (apoorvk):

Yeah, am guessing that you mean to solve this linear D.E, and find the *y* actually..

OpenStudy (anonymous):

yeah...

OpenStudy (unklerhaukus):

\[(1-x)\frac{\text dy}{\text dx}-xy=1\] \[\frac{\text dy}{\text dx}-\frac{xy}{(1-x)}=\frac 1{(1-x)}\] \[y'(x)+\frac {-x}{1-x}y=\frac 1{(1-x)}\]

OpenStudy (anonymous):

ok what's the answer? is dy/dx = (1+xy)) / (1-x) correct? my steps are simple--> divide both sides by (1-x) thus dy/dx - xy/(1-x) = 1 / (1-x) then dy/dx = 1/(1-x) + xy / (1-x) then dy/dx = (1+xy) / (1-x) [ common denominator, simple alegbra] I'm newbie to implicit differentiiation. please correct me.

OpenStudy (unklerhaukus):

\[(1-x)y'(x)-xy=1\] \[y'(x)+\frac {-x}{1-x}y=\frac 1{(1-x)}\]\[y'(x)=\frac 1{(1-x)}+\frac {x}{1-x}y\]

OpenStudy (unklerhaukus):

what you have done looks right

OpenStudy (anonymous):

thank a lot. It does not appear to involve differentiation at all.

OpenStudy (unklerhaukus):

is the question asking for \(y'\) or \(y\)

OpenStudy (anonymous):

What's the difference if it is y' or y? I am not sure? Thanks a lot.

OpenStudy (unklerhaukus):

well to you think you are being asked to isolate y', which has been done and dosent require any integration, or to find y which requires a few integrations